OpenStudy (anonymous):

to find the slope here in this problem y= -7x +3 the answer would be -10 am i right or not

OpenStudy (aroub):

no! :( / y=mx+b , m is the slope

OpenStudy (anonymous):

The slope is -7

OpenStudy (aroub):

so what do u get y=-7x+3 y=mx+b the answer is -7